OpenStudy (anonymous):

1) log x^2 with base 6 = 4 help me please 2) log x^1/2 with base 2 = 1/2 3) 8^(x^2-x) = 32 ^(1-x)

hartnn (hartnn):

so you know \(\Large \log_ab=c \implies b=a^c \) ?

OpenStudy (anonymous):

Hey, you again :)

hartnn (hartnn):

hi, yes :)

OpenStudy (anonymous):

yes. you're amazing the way you teach, youll let the viewer understand

hartnn (hartnn):

thank you for the compliment! i feel happy when i make someone understand something :)

hartnn (hartnn):

so ,what about log x^2 with base 6 = 4 ?

OpenStudy (anonymous):

6^4 = x^2?

hartnn (hartnn):

and you were typing something ? feel free to interact :)

hartnn (hartnn):

absolutely correct! just take square root on both sides to get 'x'

hartnn (hartnn):

so which 2 values of x did u get ?

OpenStudy (anonymous):

soo its 36! ^_^

OpenStudy (anonymous):

am i right?

hartnn (hartnn):

yes! and other value ?

hartnn (hartnn):

oh, wait!

OpenStudy (anonymous):

is there any other value?

hartnn (hartnn):

yes, x^2 = 36^2 ----> x^2-36^2 = 0 (x+36)(x-36) =0 x = -36, x=36 got this ?

hartnn (hartnn):

while taking square root we consider both \(\pm\) values

hartnn (hartnn):

make sense ?

OpenStudy (anonymous):

anyway why is it that 36 was raised by 2?

hartnn (hartnn):

6^4 = (6^2)^2 = (36)^2

OpenStudy (anonymous):

number 2 pleasee

myininaya (myininaya):

\[\log_2(x^2)=\frac{1}{2}\] Try write in exponential form. \[\log_a(y)=x \text{ equivalent to } a^x=y \]

OpenStudy (anonymous):

2^1/2 = x ^ 1/2

OpenStudy (anonymous):

2^1/2 = x ^ 1/2

myininaya (myininaya):

the inside of the log is x^2, correct?

myininaya (myininaya):

well for my problem it is

myininaya (myininaya):

for your problem above inside is x^1/2

myininaya (myininaya):

Ok what does that imply x is?

myininaya (myininaya):

\[x^\frac{1}{2}=2^\frac{1}{2} \] just look by inspect... what would x have to be in order for both sides to be the same.

OpenStudy (anonymous):

so x would be 2/? thanlks

myininaya (myininaya):

yes x would be 2 since 2^1/2=2^1/2

OpenStudy (anonymous):

how about my problem number 3

myininaya (myininaya):

Write 8 as 2^something Write 32 as 2^something

myininaya (myininaya):

Goal get the same base on both sides.

myininaya (myininaya):

2^1=2 2^2=2*2=4 2^3=2*2*2=4(2)=8 2^4=2*2*2*2=8(2)=16 .... What is 2^5?

OpenStudy (anonymous):

8 as 2^3 and 32 = 2 ^ 5

myininaya (myininaya):

yes

myininaya (myininaya):

\[2^{3(x^2-x)}=2^{5(1-x)}\]

myininaya (myininaya):

if a^x=a^y then a=y.

myininaya (myininaya):

So set your exponents equal and solve for x.

OpenStudy (anonymous):

so 3 (x^2 - 3x) = 5(1-x) shall i distribute it?

myininaya (myininaya):

yes

OpenStudy (anonymous):

3x^2 - 3x = 5 = 5x what would be next?

myininaya (myininaya):

3(x^2-x)=5(1-x) 3x^2-3x=5-5x Put everything on one side.

OpenStudy (anonymous):

3x^2 + 2x = 5?

myininaya (myininaya):

Ok then subtract 5 on both sides. You can then use the quadratic formula.

OpenStudy (anonymous):

you mean 3x^2 + 2x - 5 = 0 than factor it out?

myininaya (myininaya):

You can use quadratic formula or factoring.

OpenStudy (anonymous):

how is it done please

OpenStudy (anonymous):

i find it hard on factoring

myininaya (myininaya):

What two factors of a*c have product ac and sum b? What two factors of -15 have product -15 and sum 2?

OpenStudy (anonymous):

gonna finish it tomorrow, i'll study this buy tomorrow for it's time for me to sleep. it\s already 1 am here. how about there?

OpenStudy (anonymous):

thankyou so much for helping me. see ya

myininaya (myininaya):

noon time

OpenStudy (anonymous):

aw, yeah, so goodbye for bow. I'll be right back. thankyou so much! its time for me to sleep

OpenStudy (anonymous):

*now

hartnn (hartnn):

have sweet dreams! :)